OpenStudy (ibbutibbu.):

Can I please get some help? Find the point-slope form of the equation that represents the line that passes through the points (-3, 2) and (2, 1).

OpenStudy (anonymous):

First you have to determine he slop of the line using \[slope(m)=\frac{ y2-y1 }{ x2-x1 } \] after finding the slope you use the point-slope formula, which is \[y-y1=m(x-x1)\] y is your uotput and x your input just substitute and thats your answer

OpenStudy (ibbutibbu.):

Thank You @joanip_

OpenStudy (anonymous):

You're welcome :)

OpenStudy (ibbutibbu.):

Wait... What point do you use in the point slope formula...? @joanip_

OpenStudy (anonymous):

The first one, since the variables are x1 and y1

OpenStudy (ibbutibbu.):

Ok, thanks :)

OpenStudy (anonymous):

No problem